OpenStudy (anonymous):

Probability question: You have a spinner with orange, red, blue, yellow, and green sections. What are the odds against spinning a red and an orange.

OpenStudy (anonymous):

2/5

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

well you have 5 different colors on the wheel. you spin the wheel you have 2 out of 5 colors to get red or orange. so that is where I got 2/5

OpenStudy (anonymous):

But that's the PROBABILITY of getting a red or a five. I was looking to get the ODDS AGAINST.

OpenStudy (anonymous):

The odds against spinning red or orange are (3/5)/(2/5) = 3/2 sorry read ur question wrong at first

OpenStudy (anonymous):

Where did you get the 3/5 and 2/5 ?

OpenStudy (anonymous):

The odds are 3/5. There are 5 total colors each taking the same amount of space. The odds of not landing on red and orange are the odds of landing on the remaning 3. ----------------------------------- SamanthaN The odds against spinning red or orange are (3/5)/(2/5) = 3/2 sorry read ur question wrong at first This is wrong, the spinner is not spun more than once and does not eliminate the color if it is landed on. Even if it did the odds would be (3/5)/(2/5) = 6/25.

OpenStudy (anonymous):

But I thought odds against are (odds of non-preferred outcomes/odds of preferred outcomes) ?

OpenStudy (anonymous):

Got it figured out, nevermind